# Multimodal Needle in a Haystack: Benchmarking Long-Context Capability of Multimodal Large Language Models

### v1

@@ 4.4 Statistical Significance
Fig. 3 shows the results of our hypothesis test of exact accuracy (success rate) over varying sample sizes, i.e., from 100 to 1000 samples. The solid lines indicate the exact accuracy, while the shaded areas indicate the standard error. The results show that for all models, (1) the accuracy stabilizes after $500$ samples, and (2) the standard error drops significantly as sample sizes increase from $100$ to $1000$ samples. This demonstrates (1) the necessity of using larger sample size and (2) the sufficiency of using a sample size of $1000$, to achieve reliable evaluation (see the Supplement for details and more experiments on statistical significance).
@@ Appendix A Details of the MMNeedle Dataset
Resolution of Sub-Images. As discussed in Sec. 3.2 of the main paper, we find that humans and LLMs cannot effectively recognize MS COCO images with a resolution lower than $256$. Fig. 4 shows 4 random samples with $8\times 8$ stitching from our MMNeedle dataset. As demonstrated in these images, our $256\times 256$ resolution ensures a reasonable balance of input tokens and image quality. The Azure OpenAI document states that: “If an image is ambiguous or unclear, the model will do its best to interpret it. However, the results might be less accurate. A good rule of thumb is that if an average human can’t see the info in an image at the resolutions used in low/high res mode, then the model can’t either.” The Anthropic document also states that “Ensure your images are clear and not too blurry or pixelated. Claude may struggle to accurately interpret unclear or low-quality images.”
@@
Data Source. The asset we use in our paper, i,e, MS COCO 2014 dataset, is licensed under a Creative Commons Attribution 4.0 License (see the terms of use for details). This license permits the copying, redistribution, remixing, transforming, and building upon the material for any purpose, including commercial use, provided appropriate credit is given, a link to the license is provided, and any changes made are indicated. As a user of the MS COCO dataset, we acknowledge and comply with the requirements of the CC BY 4.0 license.
Ethical Considerations. Our MMNeedle dataset, created from MS COCO images, adheres to ethical guidelines and ensures that the usage of images is respectful and does not infringe on personal privacy. We ensure that MMNeedle dataset does not contain any personally identifiable information or offensive content. We bear all responsibility in case of violation of rights and confirm that we use the CC BY 4.0 data license.
@@
Prompt for Multi-Needle Evaluation. As mentioned in Sec. 3.4 of the main paper, for $K$-needle ($K>1$) evaluation, we use the following prompt for the evaluated MLLM:
@@
Given M images indexed from 1 to M, each divided into N*N sub-images,
identify the sub-images that best match the provided K captions.
Respond in the format: ‘‘index_1, row_1, column_1; ...; index_K, row_K, column_K.’’ Only provide this information.
For example, ‘‘1, 2, 3’’ indicates the sub-image in the first image, second row, and third column.
If no sub-image matches a caption, respond with ‘‘-1’’ for that caption.
@@ Appendix C Implementation Details
All the code, data, and instructions required to reproduce the main experimental results are provided in the supplementary materials, as well as at https://github.com/Wang-ML-Lab/multimodal-needle-in-a-haystack.
@@
    • Claude 3 Opus [1] is the strongest MLLM developed by Anthropic. We use the model version claude-3-opus-20240229.
    • Gemini Pro 1.0 [24] is an advanced version of Google Gemini, offering enhanced performance in multimodal tasks. We use the model version gemini-1.0-pro-vision-latest.
    • Gemini Pro 1.5 [23] is built upon Gemini Pro 1.0 with further optimizations in multimodal capability, serving as the strongest model version of Google Gemini. We use the model version gemini-1.5-pro-latest.
    • GPT-4V [3] is an extension of OpenAI’s GPT-4, equipped with vision capabilities for multimodal tasks. We use Azure OpenAI API with the model version 2024-03-01-preview.
    • GPT-4o [2] is the latest and strongest variant of OpenAI’s GPT-4. We use Azure OpenAI API with the model version 2024-05-01-preview.
@@
    • CogVLM [25] is a state-of-the-art MLLM for single-image inputs. We evaluate CogVLM-17B-base and CogVLM2-Llama-3 (the latest and strongest version).
    • Fuyu-8B [4] is a state-of-the-art, 8-billion-parameter model that excels in multimodal tasks compared to other models of similar size.
    • mPLUG-Owl-v2 [27] is an updated version of mPLUG-Owl and also a state-of-the-art MLLM.
    • InstructBLIP [7] is another state-of-the-art MLLM for single-image inputs. We evaluate InstructBLIP-Vicuna-13B and InstructBLIP-Flan-T5-XXL, which are its two strongest variants.
    • IDEFICS2 [14] is the latest version of IDEFICS and also a state-of-the-art MLLM.
    • LLaVA-Llama-3 [16] is the latest and strongest version of LLaVA [18] and also a state-of-the-art MLLM.

### v2

@@ 4.4 Statistical Significance
Fig. 3 shows the results of our hypothesis test of exact accuracy (success rate) over varying sample sizes, i.e., from 100 to 1000 samples. The solid lines indicate the exact accuracy, while the shaded areas indicate the standard error. The results show that for all models, (1) the accuracy stabilizes after $500$ samples, and (2) the standard error drops significantly as sample sizes increase from $100$ to $1000$ samples. This demonstrates (1) the necessity of using larger sample size and (2) the sufficiency of using a sample size of $1000$, to achieve reliable evaluation (see Appendix D for details and more experiments on statistical significance).
@@ 5 Conclusion
6 Ethical Considerations
Our MMNeedle dataset, created from MS COCO images, adheres to ethical guidelines and ensures that the usage of images is respectful and does not infringe on personal privacy. We ensure that MMNeedle dataset does not contain any personally identifiable information or offensive content. We bear all responsibility in case of violation of rights and confirm that we use the CC BY 4.0 data license.
Despite these precautions, there remains a risk that the benchmark’s capabilities could be misused, particularly in scenarios where models are pushed to handle extensive visual contexts that may lead to unintended inferences or biases. Additionally, the risk of hallucination in negative samples, where the model incorrectly identifies a nonexistent target, highlights the importance of responsible use and the need for thorough evaluation before deploying these models in high-stakes applications.
7 Limitations
Our MMNeedle Benchmark assumes that the evaluated MLLM can understand and follow both visual and textual instructions, and that the model can process multiple images as input in a single query. While this is not general, we note that these assumptions (and capabilities) are necessary for modern, state-of-the-art MLLMs. Adding textual or visual index labels next to each image or sub-image could potentially enhance the performance of models. However, we leave this exploration for future work for the following reasons: (1) Our MMNeedle’s goal is to measure MLLM’s long-context capability on natural images. Accuracy of predicting sub-image indices serves as one way of measuring such capabilitiy, but the accuracy itself is not the final goal. (2) This approach alters the original image content. MMNeedle is also limited by the supported number $M$ and stitching size $N$ of image inputs in MLLMs. However, our framework can seamlessly accommodate larger $M$ and $N$ once open-source and API models (e.g., GPT-4o) begin to support them.
8 Acknowledgements
We sincerely appreciate the generous support from the Microsoft Research AI & Society Fellowship, NSF Grant IIS-2127918, NSF CAREER Award IIS-2340125, NIH Grant 1R01CA297832, and the Amazon Faculty Research Award. This research is also supported by NSF National Artificial Intelligence Research Resource (NAIRR) Pilot and the Frontera supercomputer, funded by the National Science Foundation (award NSF-OAC 1818253) and hosted at the Texas Advanced Computing Center (TACC) at The University of Texas at Austin. Finally, we extend our gratitude to the Center for AI Safety (CAIS) for providing the essential computing resources that made this work possible.
@@ Appendix A Details of the MMNeedle Dataset
Limits on the Image Numbers. We set the maximum number of complete images to $M=10$ because this is the largest number of input images that GPT-4V/4o can support. Note that our framework can easily handle larger $N$ and $M$ once open-source and API models (e.g., GPT-4o) start to support them. Table 6 below summarizes each API-based model’s limit for the number of images.
Table 6: Maximum number of images per request. "*" indicates that the OpenAI GPT-4V/4o API also supports a maximum of $10$ images with high quality. Other numbers are hard limits.
API-Based Model Limit
Azure GPT-4V/4o 10
OpenAI GPT-4V/4o 10
Claude 3 Opus 20
Gemini 1.0 Pro 16
It is worth noting that:
  • Azure OpenAI API only supports $10$ images for GPT-4V/4o. For example, an Azure document states that “When uploading images, there is a limit of 10 images per chat request.” Another Azure document states that “GPT-4o max images per request” is $10$.
  • Regular OpenAI API also supports a maximum of 10 images with high quality. Specifically, an OpenAI document states that “the token cost of a given image is determined by two factors: its size, and the detail option on each image_url block”. Therefore, to ensure sufficient quality/resolution of image inputs, we cannot upload more than $10$ images to GPT-4V/4o in the MMNeedle benchmark.
  • Other models also have a limit on the number of input images (e.g., $20$ for Claude and $16$ for Gemini). Specifically, the Claude 3 Opus document states that “You can include multiple images in a single request (up to 5 for claude.ai and 20 for API requests)”, and the Gemini 1.0 Pro Vision supports up to “16 images” as “Maximum number of images per request”.
Therefore, to ensure a fair comparison, we conducted all multi-image experiments on the $M=10$ images setting.
Purpose of Image Stitching. The reason we introduce stitching with $N\times N>1\times 1$ is as follows:
  • API-based models, such as GPT-4V/4o, can support at most $10$ images as inputs, which is surprisingly small. To further evaluate long contexts with more images, we decided to introduce image stitching. As a result, when $M=10,N\times N=8\times 8$, there are equivalently $640$ sub-images in the context, which is sufficiently large compared to the API limits of a few images.
  • Image stitching enables us to conduct additional evaluation on MLLMs’ capability in localization and retrieval of sub-images within the complete input images, which is another important aspect of long-context problems.
@@
Resolution of Sub-Images. As discussed in Sec. 3.2 of the main paper, we find that humans and LLMs cannot effectively recognize MS COCO images with a resolution lower than $256$. Fig. 4 shows 4 random samples with $8\times 8$ stitching from our MMNeedle dataset. As demonstrated in these images, our $256\times 256$ resolution ensures a reasonable balance of input tokens and image quality. Consequently, for a stitch size of $N\times N$, the overall resolution becomes $256N\times 256N$, resulting in a longer input context length that scales linearly with the stitch size $N$. This approach ensures that we do not downsample the sub-images in the stitched image, while still maintaining high image quality for the model’s comprehension. The Azure OpenAI document states that: “If an image is ambiguous or unclear, the model will do its best to interpret it. However, the results might be less accurate. A good rule of thumb is that if an average human can’t see the info in an image at the resolutions used in low/high res mode, then the model can’t either.” The Anthropic document also states that “Ensure your images are clear and not too blurry or pixelated. Claude may struggle to accurately interpret unclear or low-quality images.” Indeed, our stitched images demonstrate sufficiently high resolution to be recognized by both humans and MLLMs, and there is very little content loss or noise introduced.
@@
Data Source. The asset we use in our paper, i,e, MS COCO 2014 dataset, is licensed under a Creative Commons Attribution 4.0 License. This license permits the copying, redistribution, remixing, transforming, and building upon the material for any purpose, including commercial use, provided appropriate credit is given, and any changes made are indicated. As a user of the MS COCO dataset, we acknowledge and comply with the requirements of the CC BY 4.0 license.
@@
Prompt Design For single-needle evaluation, we use the following prompt for the evaluated LLM:
[…]
        + "Caption: " + Caption
where the instructions to MLLMs is as follows:
Given $M$ images indexed from $1$ to $M$, each divided into $N\times N$ sub-images, identify the sub-image that best matches the provided caption. Respond with “index, row, column” and nothing else. For example, “1, 2, 3” indicates the sub-image in the first image, second row, and third column. If no match is found, respond only with “-1”.
We use a similar prompt for the multi-needle setting. Specifically, for $K$-needle ($K>1$) evaluation, we use the following prompt for the evaluated MLLM:
[…]
@@
Given $M$ images indexed from $1$ to $M$, each divided into $N\times N$ sub-images, identify the sub-images that best match the provided K captions. Respond in the format: “index_1, row_1, column_1; …; index_K, row_K, column_K.” Only provide this information. For example, “1, 2, 3” indicates the sub-image in the first image, second row, and third column. If no sub-image matches a caption, respond with “-1” for that caption.
@@ Appendix C Implementation Details
All the code, data, and instructions required to reproduce the main experimental results are provided in the supplementary materials (“Software” and “Data”).
@@
    • Claude 3 Opus ant (2023) is the strongest MLLM developed by Anthropic. We use the model version claude-3-opus-20240229.
    • Gemini Pro 1.0 Team et al. (2023) is an advanced version of Google Gemini, offering enhanced performance in multimodal tasks. We use the model version gemini-1.0-pro-vision-latest.
    • Gemini Pro 1.5 Reid et al. (2024) is built upon Gemini Pro 1.0 with further optimizations in multimodal capability, serving as the strongest model version of Google Gemini. We use the model version gemini-1.5-pro-latest.
    • GPT-4V Achiam et al. (2023) is an extension of OpenAI’s GPT-4, equipped with vision capabilities for multimodal tasks. We use Azure OpenAI API with the model version 2024-03-01-preview.
    • GPT-4o ope (2024) is the latest and strongest variant of OpenAI’s GPT-4. We use Azure OpenAI API with the model version 2024-05-01-preview.
@@
    • CogVLM Wang et al. (2023) is a state-of-the-art MLLM for single-image inputs. We evaluate CogVLM-17B-base and CogVLM2-Llama-3 (the latest and strongest version).
    • Fuyu-8B Bavishi et al. (2023) is a state-of-the-art, 8-billion-parameter model that excels in multimodal tasks compared to other models of similar size.
    • mPLUG-Owl-v2 Ye et al. (2023) is an updated version of mPLUG-Owl and also a state-of-the-art MLLM.
    • InstructBLIP Dai et al. (2024) is another state-of-the-art MLLM for single-image inputs. We evaluate InstructBLIP-Vicuna-13B and InstructBLIP-Flan-T5-XXL, which are its two strongest variants.
    • IDEFICS2 Laurençon et al. (2024) is the latest version of IDEFICS and also a state-of-the-art MLLM.
    • LLaVA-Llama-3 Li et al. (2024) is the latest and strongest version of LLaVA Liu et al. (2024) and also a state-of-the-art MLLM.